LASER SPECTROSCOPY FOR STABLE-ISOTOPE ANALYSIS OF ATMOSPHERIC MOLECULES

\underline{PIN CHEN}, EARTH & SPACE SCIENCES DIVISION, JET PROPULSION LABORATORY, CAL-IFORNIA INSTITUTE OF TECHNOLOGY, PASADENA, CA 91109.

Stable-isotope analysis provides valuable constraints on the global budgets of many important atmospheric species (e.g. CO_{2}, CH_{4}, CO, and N_{2}O). High-resolution laser spectroscopy promises to substantially reduce the laborious sample pretreatments often required by isotope-ratio mass spectrometry (IRMS), the incumbent state-of-the-art, as well as being capable of readily resolving isotopic species of nearly equal masses (e.g. CH_{3}D and ^{13}CH_{4}). However, achieving sensitivity levels similar to that of IRMS has been a daunting task. This talk will provide a brief review of the progress and accomplishment to date in the application of laser spectroscopy to atmospheric stable-isotope analysis.